OpenStudy (anonymous):

A=1/2(bh) solve for b

OpenStudy (lfreeman1285):

First, we should get rid of the 1/2, so we can multiply by 2. This will give you 2A=bh. Now you just divide by h, and then there's your answer. b=2A/h

OpenStudy (anonymous):

Divided each side by h and, divide by 1/2 and then simplify.\[b=\frac{2 A}{h} \]

OpenStudy (nincompoop):

\(\huge A = \frac{bh}{2} \rightarrow \color{red}{\frac{2}{h}} \times A = \frac{bh}{2} \times \color{red}{\frac{2}{h}}\)